During the time a compact disc (CD) accelerates from rest to a constant rotational speed of 477 rev/min, it accelerates with an angular acceleration of 794 rad/s2 . What is the angular displacement of the CD?

answer is pi/2 rad
Answer #1

Given: Initial angular velocity = \omega _{i} = 0.

Final angular velocity = w2T X frequency =  2T x 477 60rad/s.

angular acceleration = \alpha= 794 rad/s2.

Angular displacement is to be calculated, i.e \theta.

As the angular acceleration is constant, we can apply w2ab .

Therefore 3 = \frac{\pi }{2} rad.